AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 377     Entry time: Sat Aug 13 10:23:59 2022
Author: JM, TD 
Subject: Noise tests pt3 and pulser walk 

Waveforms and discriminator ON. Pulser re-connected to all fees via daisy chain. V=100v, I=8.36uA

 

Pulser peak widths (ch FWHM)

aida01 ~ 109

aida02 ~ 98

aida05 ~ 24

aida03 ~ 118

aida04 ~ 136

Attachments 2, 3 & 4 1.8.W spectra, 20us FSR

Attachment 5 FEE64 temps OK

Attachment 6 per FEE64 ADC data item rates OK

Attachment 7 per FEE64 rate spectra

 

Merger reset following procedure from https://elog.ph.ed.ac.uk/DESPEC/36. Keywords for future elog search NetVar, xfer links

 

 

Pulser walkthrough R94

EG&G Ortec 448 Pulser settings

Amplitudes 90,000 - 10,000 @ 10,000 step

Attenuators x5 IN

t_r 50ns tau_d 50us

polarity + (- polarity via Cooknell SA1)

Frequency int relay

Motors box turned ON, motors pi ON, dlink network connector ON. Motors box, pi and dlink all in 19 inch rack but power supply is from a uk mains extender daisy chained to vacuum mains extender.  19 inch rack power supply remains on clean line.

Pulser peak widths (ch FWHM)

aida01 ~ 105

aida02 ~ 98

aida05 ~ 23

aida03 ~ 132

aida04 ~ 144

See attachments 8 & 9

 

Detectors moved in by pnuematic only. Leakage current increased to 9.6 uA (same as leakage current during bias ramping)

Pulser Walk R95

Pulser settings as above

See attachments 10 & 11

 

Detectors moved back out . Motors box, pi and dlink still turned ON. Leakage current to 8.36 uA

Pulser Walk R96

Pulser settings as above

See attachments 1 & 12

 

 

Jumper removed from nn bias fee adaptor board (aida04). No jumpers on any adaptor boards now. V=100V, I=8.36uA

See attachments 14 & 15 - 1.8.W spectra 20us FSR

Pulser on 90,000 x5 attenuate

aida01 ~ 123, centroid ~14691

aida02 ~ 106, centroid ~ 15038

aida05 ~ 23, centroid ~ 14414

aida03 ~ 125, centroid ~ 50838

aida04 ~ 127, centroid ~ 51754

 

Jumper reinstalled on aida04 nn bias. Note only link 1, link2 inaccesible with fee inserted into ERNI. See attachment 13

aida01 ~ 123, centroid ~14688

aida02 ~ 114, centroid ~ 15034

aida05 ~ 23, centroid ~ 14411

aida03 ~ 118, centroid ~ 50854

aida04 ~ 135, centroid ~ 51697

 

Pulser off. Background run to be left overnight started 14:24 R97. V=100v, I=8.36 uA.

Attachment 1: 51.PNG  265 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
51.PNG
Attachment 2: 40.PNG  288 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
40.PNG
Attachment 3: 41.PNG  359 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
41.PNG
Attachment 4: 42.PNG  308 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
42.PNG
Attachment 5: 43.PNG  210 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
43.PNG
Attachment 6: 44.PNG  196 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
44.PNG
Attachment 7: 45.PNG  267 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
45.PNG
Attachment 8: 46.PNG  269 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
46.PNG
Attachment 9: 47.PNG  275 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
47.PNG
Attachment 10: 48.PNG  269 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
48.PNG
Attachment 11: 49.PNG  265 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
49.PNG
Attachment 12: 50.PNG  268 kB  Uploaded Sat Aug 13 13:07:13 2022  | Hide | Hide all
50.PNG
Attachment 13: 20220811_172335.jpg  1.420 MB  Uploaded Sat Aug 13 14:20:00 2022  | Hide | Hide all
20220811_172335.jpg
Attachment 14: 53.PNG  311 kB  Uploaded Sat Aug 13 20:14:01 2022  | Hide | Hide all
53.PNG
Attachment 15: 52.PNG  372 kB  Uploaded Sat Aug 13 20:14:01 2022  | Hide | Hide all
52.PNG
ELOG V3.1.4-unknown